### v2

@@ 1. Introduction
    We construct a dynamic traffic flow simulation scheme to simulate any possible emergency traffic events, greatly expanding the current research space.
@@ 3. Notation
Reward $\mathcal{R}$. Referring to the design of Alegre et al. Alegre et al. (2021), we first define the vehicle waiting time. At time $t$, the total waiting time of all vehicles stopped at intersection (agent) $i$ is denoted as $W^{t}_{i}$. Then the reward of the agent is $r^{t}_{i}=W^{t-1}_{i}-W^{t}_{i}$. Our goal is to maximize the reward, which means that the agent should try to make the current waiting time shorter than the previous waiting time. The final reward is expected to converge to around 0, i.e., the system reaches a state of equilibrium. The advantage of considering waiting time as a reward is that the agent will not deliberately delay the release time of some lanes due to fewer cars there, but instead balanced take all vehicles into consideration.
There are many existing reward functions. For example, the reward value can increase with the decrease in the number of blocked vehicles, or set a pressure indicator Wei et al. (2019a) to measure the difference between the number of vehicles entering and leaving the lane. We test various reward functions in ”ingolstadt21” Ault and Sharon (2021), and this scenario is completely different from ours. We adopt the same algorithm independent PPO (IPPO) for all experiments. In this case, we evaluate various indicators and determine that the aforementioned method is the best choice, with superior performance compared to other methods. The experimental results are shown in Table 1.
Table 1. Comparison of different reward method. The first column is the various reward targets, and the first row is the system indicators. Arrows indicate the better direction, and standard deviations in brackets are obtained from multiple experiments.
@@
Policy $\mathcal{\pi}$. The decision made by agent $i$ in time $t$ based on the current partial observation $o^{t}_{i}$ is given by policy function $\pi^{t}_{i}(a^{t}_{i}|o^{t}_{i})$. The agent policy should maximize the total reward $\sum^{T}_{t=\tau}\gamma^{t-\tau}r^{t}_{i}$, where $\gamma$ is the discount factor, usually 0.98. This means that agents discount future reward, and care first about near-term reward.
@@ 4. Methodology
In this section, we will introduce the implementation of MacLight, including information aggregation, VAE feature compression, PPO, and method of dynamic traffic flow construction.
@@ 4.1. Multi-scene aggregation matrix
Clearly, the grid-based setting is the foundation for adopting CNN as the representation model. Although real-world road networks do not appear as regular as pixel grids, considering that most intersections are four-armed, the grid-like characteristics can still be observed when transforming them into a graph.
@@ 4.4. Dynamic traffic flow construction
In current TSC RL studies, the deployment of traffic flow is typically fixed, with all vehicles following predetermined routes. Our experiment, however, is the first to build a dynamic traffic flow environment that simulates emergency road events, which cause sudden changes in the distribution of traffic on other roads.
As illustrated in Fig. 3, when four central roads—D3C3, D3D2, D2C2, and C3C2—are blocked, all vehicles are asked to reroute. We conduct two experiments to analyze the effects. In Fig. 4, the gray curve shows the traffic flow distribution on a specific lane under normal conditions (without any interference), while the red curve shows the distribution after blocking a specific road. The lane blockage period is marked by a blue vertical dashed line. During this time, the traffic volume on the affected lanes drops significantly as vehicles select new optimal routes. Fig. 5 illustrates the changes in traffic flow on other, unblocked roads. Detailed distribution statistics can be found in Appendix A.
The impact of congestion on one road can be quite complex and influence other roads in unpredictable ways. Fig. 5 shows the flow distribution for unblocked lanes after the designated roads are closed. For example, while the traffic flow distribution on roads E1E10 remains largely unaffected, there is a significant increase in traffic on D4C4. On the other hand, B1A1 sees a sharp rise in traffic, while B3B4 experiences a decrease. Such interdependencies are difficult to model accurately but are common in real-world traffic systems. Consequently, our algorithm must account for these complex interactions.

### v3

@@ 1. Introduction
    We construct a dynamic traffic flow simulation scheme to simulate any possible emergency traffic event, greatly expanding the current research space.
@@ 3. Notation
Reward $\mathcal{R}$. Referring to the design of Alegre et al. Alegre et al. (2021), we first define the waiting time of the vehicle. At time $t$, the total waiting time of all vehicles stopped at intersection (agent) $i$ is denoted as $W^{t}_{i}$. Then the agent’s reward is $r^{t}_{i}=W^{t-1}_{i}-W^{t}_{i}$. Our goal is to maximize the reward, which means that the agent should try to make the current waiting time shorter than the previous waiting time. The final reward is expected to converge to around 0, that is, the system reaches a state of equilibrium. The advantage of considering waiting time as a reward is that the agent will not deliberately delay the release time of some lanes due to fewer cars there, but instead balanced take all vehicles into consideration.
There are many reward functions. For example, the reward value can increase with the decrease in the number of blocked vehicles or set a pressure indicator Wei et al. (2019a) to measure the difference between the number of vehicles entering and leaving the lane. We test various reward functions in ”ingolstadt21” Ault and Sharon (2021), and this scenario is completely different from ours. We adopt the same independent PPO (IPPO) for all experiments. In this case, we evaluate various indicators and determine that the aforementioned method is the best choice, with superior performance compared to other methods. The experimental results are shown in Table 1.
Table 1. Comparison of different reward methods. The first column includes various reward targets and a baseline, and the first row is the system indicators. Arrows indicate the better direction, and standard deviations in brackets are obtained from multiple experiments. The fixed time is similar to the real-life solution, that is, the fixed time switching signal, which is a baseline.
@@
Policy $\mathcal{\pi}$. The decision made by agent $i$ in time $t$ based on the current partial observation $o^{t}_{i}$ is given by the policy function $\pi^{t}_{i}(a^{t}_{i}|o^{t}_{i})$. The agent policy should maximize the total reward $\sum^{T}_{t=\tau}\gamma^{t-\tau}r^{t}_{i}$, where $\gamma$ is the discount factor, usually 0.98. This means that agents discount future reward and care first about near-term reward.
@@ 4. Methodology
In this section, we will introduce the implementation of MacLight, including information aggregation, VAE feature compression, PPO and method of dynamic traffic flow construction.
@@ 4.1. Multi-scene aggregation matrix
Clearly, the grid-based setting is the basis for adopting CNN as a representation model. Although real-world road networks do not appear as regular as pixel grids, considering that most intersections are four-armed, the grid-like characteristics can still be observed when transforming them into a graph.
@@ 4.4. Dynamic traffic flow construction
In current TSC RL studies, the deployment of traffic flow is typically fixed, with all vehicles following predetermined routes. However, our experiment is the first to build a dynamic traffic flow environment that simulates emergency road events, which cause sudden changes in the distribution of traffic on other roads.
As illustrated in Fig. 3, when four central roads—D3C3, D3D2, D2C2 and C3C2—are blocked, all vehicles are asked to reroute. We conduct two experiments to analyze the effects. In Fig. 4, the gray curve shows the traffic flow distribution on a specific lane under normal conditions (without interference), while the red curve shows the distribution after blocking a specific road. The lane blockage period is marked by a blue vertical dashed line. During this time, the traffic volume on the affected lanes drops significantly as vehicles select new optimal routes. Fig. 5 illustrates the changes in traffic flow on other unblocked roads. Detailed distribution statistics can be found in Appendix A.
The impact of congestion on one road can be quite complex and influence other roads in unpredictable ways. Fig. 5 shows the flow distribution for unblocked lanes after designated roads are closed. For example, while the traffic flow distribution on roads E1E10 remains largely unaffected, there is a significant increase in traffic on D4C4. On the other hand, B1A1 sees a sharp rise in traffic, while B3B4 experiences a decrease. Such interdependencies are difficult to model accurately but are common in real-world traffic systems. Consequently, our algorithm must account for these complex interactions.